OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (anonymous):

Dosent have to be long.

OpenStudy (anonymous):

It may not have to be long, but I'm sure it has to be your own words. Do you have a question about solving this system of equations?

OpenStudy (usarmy3947):

ok i have the answer but there will be a medal right?

OpenStudy (anonymous):

yes there will be

OpenStudy (usarmy3947):

x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10 : Look at the equations, see if we can find a multiplier for one of the equations that will eliminate two variables at the same time : There is! Multiply the last equation by -1 and add all three equations x – 5y + 2z = 0 x + 4y – z = 12 -2x +y -3z = -10 ----------------------Addition eliminates x and y, find z 0 + 0 - 2z = 2 Multiply by -1 2z = -2 divide both sides by 2 z = -1; they only asked for 1 variable

OpenStudy (usarmy3947):

did this help?

OpenStudy (usarmy3947):

thx you nice helping u puertoricangirl96

OpenStudy (usarmy3947):

is there another question?

OpenStudy (anonymous):

no.

OpenStudy (usarmy3947):

oh ok thx anyway if i can be of any further assistance just message me

OpenStudy (usarmy3947):

Que tengan un buen día señorita.